Deidre is picking out some movies to rent, and she is primarily interested in mysteries and action movies. She has narrowed down her selections to 10 mysteries and 18 action movies. Step 1 of 2 : How many different combinations of 4 movies can she rent?
Answer #1

10 mysteries and 18 action movies

How many different combinations of 4 movies can she rent?

0 mysteries and 4 action movies

10C0*18C4 = 2856

1 mysteries and 3 action movies

10C1*18C3 = 8160

2 mysteries and 2 action movies

10C2*18C2 =  6885

3 mysteries and 1 action movies

10C3*18C1 = 1200

4 mysteries and 0 action movies

10C4*18C0 = 210

different combinations she can rent = 2856 + 8160 + 6885 + 1200 + 210

different combinations she can rent = 19311